OpenStudy (anonymous):

If x+3/3=y+2/2, then x/3=___?

OpenStudy (whpalmer4):

Is that \[\frac{x+3}3 = \frac{y+2}2\]?

OpenStudy (anonymous):

x+3/3=y+2/2= 3(y+2)=2(x+3) 3y+6=2x+6 -6 -6 3y=2x

OpenStudy (whpalmer4):

okay, so x/3 =?

OpenStudy (anonymous):

2y?

OpenStudy (whpalmer4):

No \[3y = 2x\]Divide both sides by 2\[\frac{3y}{2} = x\]What is \(x/3 =\)?

OpenStudy (anonymous):

I don't really understand what I am supposed to do, 3y=2x /2 /2 3/2y=x

OpenStudy (whpalmer4):

If \[x=3y/2\]then what does \[x/3=\]?

OpenStudy (whpalmer4):

What is (3y/2)/3?

OpenStudy (anonymous):

.5

OpenStudy (whpalmer4):

No!!! \[x=\frac{3y}2\]divide both sides by 3\[\frac{x}3=\frac{3y}{2*3}=\]

OpenStudy (anonymous):

could you explain, im honestly so clueless right now.

OpenStudy (whpalmer4):

Simplify the fraction on the right

OpenStudy (anonymous):

3y/6 which can be to 1/2

OpenStudy (whpalmer4):

Where did y go?!?

OpenStudy (anonymous):

so 1/2y?

OpenStudy (whpalmer4):

\(\frac{1}2y\) or \(\frac{y}2\)